# Terrestrial Matter Effects on Reactor Antineutrino Oscillations: Constant vs. Fluctuated Density Profiles

### v1

@@ Abstract
The JUNO Collaboration has recently released its first reactor antineutrino oscillation result, achieving unprecedented precision in the measurement of $\Delta m^{2}_{21}$ and $\sin^{2}\theta_{12}$. We emphasize that the accurate determination and modeling of the terrestrial matter density profile are fundamental for extracting the oscillation parameters and probing the neutrino mass ordering. This paper presents a realistic piecewise-constant model for the shallow crustal density profile along the baselines from Taishan and Yangjiang to the experimental hall, based on geological and petrophysical information. The uncertainty in the density profiles arises from variations in the density and length of each segment, both of which are conservatively estimated to be 10%. A careful comparison of constant and fluctuated density profiles is provided and the implications for the precision measurement of oscillation parameters are discussed. Finally, we also discuss the possibility of tomography of the shallow Earth’s crust in future reactor neutrino experiments.
@@ 1 Introduction
The global neutrino program in the coming decades is organized around these promising frontiers. Next-generation long-baseline accelerator experiments such as Hyper-Kamiokande [14] and DUNE [15] aim to precisely measure the atmospheric oscillation parameters and the CP-violating phase $\delta$. On the reactor side, the Jiangmen Underground Neutrino Observatory (JUNO) [16, 17] is a next-generation oscillation experiments aiming at the determination of neutrino mass ordering [18, 19] and measuring oscillation parameters with sub-percent precision [20]. All of these next-generation neutrino oscillation experiments will operate in an era of high precision and require better control of systematic uncertainties on both the theoretical and experimental sides.
The JUNO experiment, the first of these next-generation flagship experiments to begin operations, has published its initial reactor antineutrino oscillation results [21, 22]. It achieved unprecedented precision in measuring $\Delta m^{2}_{21}$ and $\sin^{2}\theta_{12}$, with uncertainties of 1.55% and 2.81%, respectively. With this accuracy level, the corrections due to terrestrial matter effects [23, 24] on $\Delta m^{2}_{21}$ and $\sin^{2}\theta_{12}$ will be at the percent level [25, 26], and cannot be neglected in the analysis. In the previous analyses, a constant density profile, characterized by an average crustal density, was assumed and deemed reasonably accurate. However, since the majority of the neutrino propagation trajectories are relatively shallow, the density variation may be significant depending on the local geological conditions. Therefore, it is important to evaluate the effect of a variable density profile on the survival probability and its potential impact on the oscillation parameters.
In this work, we focus on the impact of realistic crustal density variations along the JUNO baselines. By utilizing geological maps and petrophysical data from South China, we construct a piecewise-constant matter density model for the density profiles extending from the Taishan and Yangjiang nuclear power plants to the experimental hall. We evaluate the uncertainties in the density profiles by incorporating detailed local information regarding crustal density and geometric baseline variations in each segment. Using an efficient semi-analytical method based on the Cayley-Hamilton theorem, we compare the differences in neutrino survival probabilities between the variable density profile and its constant density approximation, and discuss the potential impact on future measurements of oscillation parameters. Finally, we explore the possibility of shallow Earth crust tomography in future reactor neutrino experiments.
@@ 2.1 Geological model and baseline discretizations
The JUNO detector is located about $650~\mathrm{m}$ underground, whereas the reactors are slightly above sea level. Reactor antineutrinos thus propagate predominantly through the upper crust, over relatively shallow depths. In this regime, lateral variations of the density are dominated by changes in outcropping strata and lithology rather than by large-scale crust–mantle structures. Under these conditions, it is reasonable to approximate each surface geological unit by a representative bulk density; residual small-scale variations are expected at the level of a few percent. In the following sections we adopt segment-wise fluctuations up to $\pm 10\%$ as a deliberately conservative bound that comfortably covers these geophysical uncertainties.
@@
Figure 1: Topographic (left) and geological (right) maps around the JUNO detector site and the Taishan (TS) and Yangjiang (YJ) reactor complexes. The white lines indicate the baselines from the reactor cores to JUNO.
@@ 2.2 Fluctuations of the matter density profile
Figure 2: Geological sampling from Taishan to Jiangmen; the upper panel shows the surface elevation, and the lower panel shows the matter density of surface rocks along the TS–JUNO baseline.
Figure 3: Geological sampling from Yangjiang to Jiangmen; the upper panel shows the surface elevation, and the lower panel shows the matter density of surface rocks along the YJ–JUNO baseline.
@@ 3 Neutrino oscillations in matter
$P(\overline{\nu}_{e}\rightarrow\overline{\nu}_{e})=1$ $-\sin^{2}2\theta_{12}\cos^{4}\theta_{13}\sin^{2}\frac{\Delta m_{21}^{2}L}{4E_{\nu}}$ (3.3)
@@ 4 Numerical results
First, we examine the absolute magnitude of variable matter effects on the $\bar{\nu}_{e}\to\bar{\nu}_{e}$ survival probability. As shown in Fig. 4, the absolute differences between vacuum survival probability $P^{\rm Vac}_{ee}$ and variable-matter survival probability $P^{\rm Vari}_{ee}$ are presented for the reference matter density profiles along the TS–JUNO and YJ–JUNO baselines. One can observe that they are at the percent level and will be comparable to the expected uncertainties of $\Delta m^{2}_{21}$ and $\sin^{2}\theta_{12}$. This conclusion is rather consistent with Ref. [25] using the constant density approximation.
Figure 4: The absolute differences between $P_{ee}$ in matter and in vacuum using the reference matter density profiles.
@@
All these sampled representative profiles satisfy $|\delta\rho_{i}|\leq 10\%$ and $|\delta L_{i}|/L_{i}^{(0)}\leq 10\%$ for every segment, and their baseline-averaged densities lie within the $\sim 1$$3\%$ band around the reference values quantified in Sec. 2.2.
@@
Figure 5: Variable matter effects on the $\bar{\nu}_{e}\to\bar{\nu}_{e}$ survival probability along the TS–JUNO (upper panel) and YJ–JUNO (lower panel) baselines for NMO (left panel) and IMO (right panel).
Now we discuss how the variable matter effects affects the measurement of oscillation parameters and mass ordering. From the first reactor antineutrino oscillation result of JUNO [21], the precision of $\Delta m_{21}^{2}$ and $\sin^{2}\theta_{12}$ reached 1.55% and 2.81% respectively. Meanwhile, it will reach to 0.3% and 0.5% in 6 yrs and even smaller for the entire lifetime of 20 yrs.
In Fig. 6, the comparison of the survival probability differences between the oscillation parameter variation and the density profile variation is illustrated. As an illustration, Fig. 6 shows the absolute differences in the survival probability, $\big|P_{ee}^{X}-P_{ee}^{\rm BF}\big|$ for the TS–JUNO baseline and NMO, where $P_{ee}^{\rm BF}$ is the probability calculated with the best-fit values, and $P_{ee}^{X}$ with one of the parameters shifted according to the expected JUNO precision. For clarity, in each panel only one parameter is varied at a time, while the others are fixed at their best-fit values. The black and grey lines are that of variable density profile with the uncertainty of Scenario A. From the figures, one can read that the uncertainty of matter density profiles is more than one order of magnitude smaller than the current precision of oscillation parameters, but will increase to 20% or 50% for longer running time. Therefore a careful investigation and modelling of the density profiles is needed in the future.
@@
and illustrate the absolute differences of these quantities in the lower panels of Fig. 7. Note that the black and gray curves represent Scenario A, while the colored curves correspond to Scenario B. By comparing the magnitudes of these differences, we conclude that variations in matter density profiles contribute only 1% to the mass ordering sensitivity, making them relatively negligible. Therefore, we can assert that the additional uncertainty introduced by the imperfect knowledge of the matter density profile is insignificant, and the determination of mass ordering remains robust against realistic variable matter density effects.
@@ 5 Impact of anomalous density perturbations
For definiteness, let us consider a single localized anomaly with length $\Delta L_{a}$ and density $\rho_{a}$ embedded in an otherwise standard rock profile characterized by an average density $\bar{\rho}_{0}$. If the baseline length is $L$, the new baseline-averaged density becomes
Figure 6: Impact of the uncertainties of $\Delta m_{21}^{2}$, and $\theta_{12}$ on the $\bar{\nu}_{e}\to\bar{\nu}_{e}$ survival probability along the TS–JUNO (YJ–JUNO) baseline for normal mass ordering. The three panels in each column correspond to parameter uncertainties after 60 days , 6 years (6 yrs), 20 years (20 yrs) of data taking. In each panel, only one parameter is varied at a time, while the others are fixed to their best-fit values. And the black/gray lines denote $10\%$ variation of the benchmark constant density, namely, $\rho^{\rm Average}_{\rm TS}$ and $\rho^{\rm Average}_{\rm YJ}$ as shown in Figure 5.
Figure 7: Variable matter effects on the mass-ordering determination. Top row: absolute difference $P_{ee}^{\rm C,MO}$ between the constant-density survival probabilities in NMO and IMO for the TS–JUNO (left) and YJ–JUNO (right) baselines. Middle row: deviation $|P_{ee}^{\rm Vari,MO}-P_{ee}^{\rm C,MO}|$ induced by density fluctuations. Bottom row: relative deviation $|P_{ee}^{\rm Vari,MO}-P_{ee}^{\rm C,MO}|/P_{ee}^{\rm C,MO}$.
@@
Figure 8: Absolute difference $|P_{ee}^{\rm Anomaly}-P_{ee}^{\rm C}|$ as a function of antineutrino energy $E_{\nu}$ for the TS–JUNO baseline in NMO. The constant-density reference probability $P_{ee}^{\rm C}$ is computed for a uniform matter density $\rho^{\rm Average}_{\rm TS}$, equal to the baseline-averaged density defined in Eqs. (4.14) and (4.15). The three colored curves correspond to profiles in which an additional $1~\text{km}$ air segment ($\rho\simeq 0$) is inserted at the beginning (red), in the middle (blue), or at the end (green) of the baseline, while the remaining path is kept at density $\rho^{\rm Average}_{\rm TS}$. The differences stay at the level of a few $\times 10^{-4}$ or smaller over the entire JUNO energy range, illustrating the negligible impact of such localized low-density anomalies on the reactor antineutrino survival probability.
@@ 6 Summary
In this work, we have investigated the impact of realistic matter density variations on reactor antineutrino oscillations. Using geological information along the baselines from reactor nuclear power plants to the experimental hall, we constructed a piecewise-constant density model that approximates the South China crustal structure. We then allowed for conservative fluctuations of $\pm 10\%$ in the density of each segment and $\pm 10\%$ in the segment lengths, while keeping the total baseline lengths fixed. The resulting average matter densities and the conservative uncertainties are given as
@@
  • The absolute difference $|P_{ee}^{\rm C}-P_{ee}^{\rm Vari}|$ induced by realistic density variations is approximately $10^{-3}$ for the correlated density fluctuations of Scenario A, accounting for about 10% of the total matter effects. While this effect is negligible at the current precision, it should be considered and addressed for future measurements aiming at millesimal precision.
@@
From a broader perspective, the commencement of JUNO’s data collection and the release of the first reactor antineutrino oscillation results signify an important advancement in the global neutrino program. Alongside Hyper-Kamiokande and DUNE, JUNO will contribute to completing the understanding of three-flavor oscillations by determining the mass ordering and significantly enhancing the precision of the mixing angles and mass-squared differences. With the advent of the era of precision measurements in neutrino physics, the contributions of sub-leading and subtle effects may need to be reconsidered, such as the variable matter density profile considered in this work. Other such efforts include studies of indirect unitarity violation effects [48], radiative corrections to neutrino propagation in matter and to elastic neutrino-electron scattering [49, 50, 51]. Systematically incorporating these effects will further increase the reliability of oscillation experiments, enhance their discovery potential [52], and support precise measurements of neutrino oscillation parameters and searches for related new physics.

### v3

@@ Abstract
The JUNO Collaboration has recently released its first reactor antineutrino oscillation result, achieving unprecedented precision in the measurement of $\Delta m^{2}_{21}$ and $\sin^{2}\theta_{12}$. We emphasize that the accurate determination and modeling of the terrestrial matter density profile are fundamental for extracting the oscillation parameters and probing the neutrino mass ordering. This paper presents a realistic piecewise-constant model for the shallow crustal density profile along the baselines from Taishan and Yangjiang to the experimental hall, based on geological and petrophysical information. The uncertainty in the density profiles arises from variations in the density and length of each segment, both of which are conservatively estimated to be 10%. A careful comparison of constant and fluctuated density profiles is provided and the implications for the precision measurement of oscillation parameters are discussed. Finally, we also discuss the prospect of shallow crust tomography in future reactor neutrino experiments.
@@ 1 Introduction
The global neutrino program in the coming decades is organized around these promising frontiers. Next-generation long-baseline accelerator experiments such as Hyper-Kamiokande [1] and DUNE [3] aim to precisely measure the atmospheric oscillation parameters and the CP-violating phase $\delta$. On the reactor side, the Jiangmen Underground Neutrino Observatory (JUNO) [13, 8] is a next-generation oscillation experiment aiming at determining the neutrino mass ordering [26, 7] and measuring oscillation parameters with sub-percent precision [4]. All of these next-generation neutrino oscillation experiments will operate in an era of high precision and require better control of systematic uncertainties on both the theoretical and experimental sides.
The JUNO experiment, the first of these next-generation flagship experiments to begin operations, has published its initial reactor antineutrino oscillation results [5, 6]. It achieved unprecedented precision in measuring $\Delta m^{2}_{21}$ and $\sin^{2}\theta_{12}$, with uncertainties of 1.55% and 2.81%, respectively. With this accuracy level, the corrections due to terrestrial matter effects [47, 34] on $\Delta m^{2}_{21}$ and $\sin^{2}\theta_{12}$ will be at the percent level [28, 15], and cannot be neglected in the analysis. In previous analyses, a constant density profile, characterized by an average crustal density, was assumed and deemed reasonably accurate. However, since the majority of the neutrino propagation trajectories are relatively shallow, the density variation may be significant depending on the local geological conditions. Therefore, it is important to evaluate the effect of a variable density profile on the survival probability and its potential impact on the oscillation parameters.
In this work, we focus on the impact of realistic crustal density variations along the JUNO baselines. By utilizing geological maps and petrophysical data from South China, we construct a piecewise-constant matter density model along the baselines extending from the Taishan and Yangjiang nuclear power plants to the experimental hall. We evaluate the uncertainties in the density profiles by incorporating detailed local information regarding crustal density and geometric baseline variations in each segment. Using an efficient semi-analytical method based on the Cayley-Hamilton theorem, we compare the differences in neutrino survival probabilities between the variable density profile and its constant density approximation, and discuss the potential impact on future measurements of oscillation parameters. Finally, we explore the possibility of shallow Earth crust tomography in future reactor neutrino experiments.
@@ 2.1 Geological model and baseline discretizations
The JUNO detector is located about $750~\mathrm{m}$ underground, whereas the reactors are slightly above sea level. Reactor antineutrinos thus propagate predominantly through the upper crust, over relatively shallow depths. In this regime, lateral variations of the density are dominated by changes in outcropping strata and lithology rather than by large-scale crust–mantle structures. Under these conditions, it is reasonable to approximate each surface geological unit by a representative bulk density; residual small-scale variations are expected at the level of a few percent. In the following sections we adopt segment-wise fluctuations up to $\pm 10\%$ as a deliberately conservative bound that comfortably covers these geophysical uncertainties.
@@
Figure 1: Topographic (left) and geological (right) maps of the terrain and crustal structure traversed by the reactor antineutrino oscillation baselines from the Taishan (TS) and Yangjiang (YJ) nuclear power plants to the JUNO detector. The white lines indicate the TS–JUNO and YJ–JUNO baselines.
@@ 2.2 Fluctuations of the matter density profile
Figure 2: Geological sampling along the TS–JUNO baseline. The upper panel shows the surface elevation as a function of distance, and the lower panel shows the corresponding matter density of surface rocks.
Figure 3: Geological sampling along the YJ–JUNO baseline. The upper panel shows the surface elevation as a function of distance, and the lower panel shows the corresponding matter density of surface rocks.
@@ 3 Neutrino oscillations in matter
$P(\bar{\nu}_{e}\rightarrow\bar{\nu}_{e})=1$ $-\sin^{2}2\theta_{12}\cos^{4}\theta_{13}\sin^{2}\frac{\Delta m_{21}^{2}L}{4E_{\nu}}$ (3.3)
@@ 4 Numerical results
Figure 4: Top row: the electron antineutrino survival probability in matter, $P^{\rm Matter}_{ee}$, as a function of $E_{\nu}$ for the TS–JUNO (left) and YJ–JUNO (right) baselines. Bottom row: the absolute matter-induced correction with respect to vacuum, $|P^{\rm Matter}_{ee}-P^{\rm Vac}_{ee}|$. Solid and dashed curves correspond to normal (NMO) and inverted (IMO) mass ordering, respectively. The survival probabilities are calculated using the current best-fit values of the oscillation parameters [19] and the piecewise-constant matter-density profiles obtained from the geological sampling along the baselines (see Figs. 2 and 3 and Tables 1 and 2).
First, we examine the absolute magnitude of variable matter effects on the $\bar{\nu}_{e}\to\bar{\nu}_{e}$ survival probability. As shown in Fig. 4, the upper panels display the variable-matter survival probability $P^{\rm Matter}_{ee}$ as a function of $E_{\nu}$ for the reference density profiles along the TS–JUNO (left) and YJ–JUNO (right) baselines, while the lower panels show the absolute matter-induced correction with respect to vacuum, $|P^{\rm Matter}_{ee}-P^{\rm Vac}_{ee}|$. One can see that $P^{\rm Matter}_{ee}$ itself is typically $\mathcal{O}(10^{-1}\!-\!1)$ over the JUNO energy window, whereas the correction $|P^{\rm Matter}_{ee}-P^{\rm Vac}_{ee}|$ stays at the subpercent level in absolute size, corresponding to a relative spectral distortion at the $\mathcal{O}(1\%)$ level. Such a size is comparable to the expected uncertainties in extracting $\Delta m^{2}_{21}$ and $\sin^{2}\theta_{12}$, and is also consistent with Ref. [28] based on the constant-density approximation.
@@
All these sampled representative profiles satisfy $|\delta\rho_{i}|\leq 10\%$ and $|\delta L_{i}|/L_{i}^{(0)}\leq 10\%$ for every segment, and their baseline-averaged densities lie within a few-percent band around the reference values, consistent with the quantitative estimates in Sec. 2.2.
@@
Now we discuss how the matter-density variations affect the measurement of oscillation parameters and mass ordering. From the first reactor antineutrino oscillation result of JUNO [5], the precisions of $\Delta m_{21}^{2}$ and $\sin^{2}\theta_{12}$ reached 1.55% and 2.81% respectively. With more data, these uncertainties are expected to be reduced to about 0.3% and 0.5% after 6 years of running, and to even smaller values over the full 20-year lifetime.
In Fig. 6, the comparison of the survival probability differences between the oscillation parameter variation and the density profile variation is illustrated. As an illustration, Fig. 6 shows the absolute differences in the survival probability, $\big|P_{ee}^{X}-P_{ee}^{\rm BF}\big|$ for the TS–JUNO baseline and NMO, where $P_{ee}^{\rm BF}$ is the probability calculated with the best-fit values, and $P_{ee}^{X}$ with one of the parameters shifted according to the expected JUNO precision. For clarity, in each panel only one parameter is varied at a time, while the others are fixed at their best-fit values. The black and gray lines are that of variable density profile with the uncertainty of Scenario A. From the figures, one can read that the uncertainty of matter density profiles is currently more than one order of magnitude smaller than the uncertainty due to the oscillation parameters, but will grow to about $20\%$ (after 6 years) and $50\%$ (after 20 years) of the parameter-induced effect as the statistical precision improves. Therefore a careful investigation and modeling of the density profiles is needed in the future.
@@
Figure 5: Variable matter effects on the $\bar{\nu}_{e}\to\bar{\nu}_{e}$ survival probability along the TS–JUNO (upper row) and YJ–JUNO (lower row) baselines for normal (NMO, left column) and inverted (IMO, right column) mass ordering. Shown is the absolute difference between the probabilities computed with the benchmark constant-density profile, whose baseline-averaged densities $\rho^{\rm Average}_{\rm TS}$ and $\rho^{\rm Average}_{\rm YJ}$ are defined in Eqs. (4.14) and (4.15), and with variable matter density, $|P^{\rm C}_{ee}-P^{\rm Vari}_{ee}|$. The black and gray curves correspond to Scenario A, where all rock densities are coherently rescaled by $\pm 10\%$, while the colored curves show five representative realizations of Scenario B with independent $\pm 10\%$ fluctuations in each rock segment.
and illustrate the absolute differences of these quantities in the lower panels of Fig. 7. Note that the black and gray curves represent Scenario A, while the colored curves correspond to Scenario B. By comparing the magnitudes of these differences, we conclude that variations in matter density profiles contribute only $1\%$ to the mass ordering sensitivity, making them relatively negligible. Therefore, we can assert that the additional uncertainty introduced by the imperfect knowledge of the matter density profile is insignificant, and the determination of mass ordering remains robust against realistic variable matter density effects.
To put the matter-induced corrections discussed above into perspective, it is useful to compare them with other uncertainties relevant for reactor antineutrino measurements. For JUNO, detector- and reactor-related uncertainties can affect the predicted antineutrino rate at the $\sim 1\%$ level [5]. On the other hand, the inverse beta-decay (IBD) cross section can be evaluated with sub-percent (typically per-mille) accuracy in state-of-the-art treatments [41]. In comparison, the overall matter-induced modification of the survival spectrum shown in Fig. 4 is at the percent level in absolute size, while the additional impact from replacing the constant-density approximation by realistic, density profiles affects the mass-ordering discriminator only at the much smaller level. Therefore, density-profile variations constitute a subleading effect with respect to the dominant experimental systematics, but they remain relevant for future analyses aiming at per-mille precision.
@@ 5 Impact of anomalous density perturbations
For definiteness, let us consider a single localized anomaly with length $\Delta L_{a}$ and density $\rho_{a}$ embedded in an otherwise standard rock profile characterized by an average density $\bar{\rho}_{0}$.
If the baseline length is $L$, the new baseline-averaged density becomes
@@
Figure 6: Impact of the uncertainties in $\Delta m^{2}_{21}$ and $\theta_{12}$ on the $\bar{\nu}_{e}\to\bar{\nu}_{e}$ survival probability along the TS–JUNO (left column) and YJ–JUNO (right column) baselines for normal mass ordering. From top to bottom, the panels correspond to the expected parameter uncertainties after 60 days, 6 years (6 yrs), and 20 years (20 yrs) of data taking. In each panel, only one parameter is varied within its projected $1\sigma$ range, while the others are fixed at their best-fit values. The black and gray lines indicate the effect of a $\pm 10\%$ variation of the benchmark constant density, $\rho^{\rm Average}_{\rm TS}$ or $\rho^{\rm Average}_{\rm YJ}$ (Scenario A), as in Fig. 5.
Figure 7: Variable matter effects on the determination of the neutrino mass ordering. Top row: absolute difference $P^{\rm C,MO}_{ee}\equiv|P^{\rm C,NMO}_{ee}-P^{\rm C,IMO}_{ee}|$ between the constant-density survival probabilities in the normal mass ordering (NMO) and inverted mass ordering (IMO) for the TS–JUNO (left) and YJ–JUNO (right) baselines. Bottom row: deviation $|P^{\rm Vari,MO}_{ee}-P^{\rm C,MO}_{ee}|$ induced by matter-density variations. In all panels, the oscillation probabilities are computed using the current best-fit values of the oscillation parameters. The constant-density probabilities $P^{\rm C}_{ee}$ are obtained with the benchmark baseline-averaged densities $\rho^{\rm Average}_{\rm TS}$ and $\rho^{\rm Average}_{\rm YJ}$ defined in Eqs. (4.14) and (4.15). The black and gray curves correspond to Scenario A, while the colored curves represent five representative realizations of Scenario B.
@@
Figure 8: Absolute difference $|P_{ee}^{\rm Anomaly}-P_{ee}^{\rm C}|$ as a function of antineutrino energy $E_{\nu}$ for the TS–JUNO baseline in NMO. The constant-density reference probability $P_{ee}^{\rm C}$ is computed for a uniform matter density $\rho^{\rm Average}_{\rm TS}$, equal to the baseline-averaged density defined in Eq. (4.14). The three colored curves correspond to profiles in which an additional $1~\text{km}$ air segment ($\rho\simeq 0$) is inserted at the beginning (red), in the middle (blue), or at the end (green) of the baseline, while the remaining path is kept at density $\rho^{\rm Average}_{\rm TS}$. The differences stay at the level of a few $\times 10^{-4}$ or smaller over the entire JUNO energy range, illustrating the negligible impact of such localized low-density anomalies on the reactor antineutrino survival probability.
@@ 6 Summary
In this work, we have investigated the impact of realistic matter density variations on reactor antineutrino oscillations. Using geological information along the baselines from the Taishan and Yangjiang nuclear power plants to the experimental hall, we constructed a piecewise-constant density model that approximates the South China crustal structure. We then allowed for conservative fluctuations of $\pm 10\%$ in the density of each segment and $\pm 10\%$ in the segment lengths, while keeping the total baseline lengths fixed. The resulting average matter densities and the conservative uncertainties are given as
@@
  • The absolute difference $|P_{ee}^{\rm C}-P_{ee}^{\rm Vari}|$ induced by realistic density variations is approximately $10^{-3}$ for the correlated density fluctuations of Scenario A, accounting for about 10% of the total matter effects. While this effect is negligible at the current precision, it should be considered and addressed for future measurements aiming at per-mille $(10^{-3})$ precision.
@@
From a broader perspective, the commencement of JUNO’s data collection and the release of the first reactor antineutrino oscillation results signify an important advancement in the global neutrino program. Alongside Hyper-Kamiokande and DUNE, JUNO will contribute to completing the understanding of three-flavor oscillations by determining the mass ordering and significantly enhancing the precision of the mixing angles and mass-squared differences. With the advent of the era of precision measurements in neutrino physics, the contributions of sub-leading and subtle effects may need to be reconsidered, such as the variable matter density profile considered in this work. Other such efforts include studies of indirect unitarity violation effects [29], radiative corrections to neutrino propagation in matter and to elastic neutrino-electron scattering [23, 24, 21, 22]. Systematically incorporating these effects will further increase the reliability of oscillation experiments, enhance their discovery potential [42, 46], and support precise measurements of neutrino oscillation parameters and searches for related new physics.